A laboratory supervision in Type III hospital is investigating a number of reported on-the-job training accidents related to needle stick injuries over a period of one month. Based on the past records, she has the following records on the needle stick injuries that were recorded:


Needlestick injuries reported: 0, 1, 2, 3, 4, 5

Frequency: 8, 10, 2, 5, 3, 7


1
Expert's answer
2022-01-31T12:30:36-0500

Since the question is incomplete, let us assume that we are asked to find mean of this data.

Solution:

Needlestick injuries reported "(x)" : 0, 1, 2, 3, 4, 5

Frequency "(f)" : 8, 10, 2, 5, 3, 7

Mean"=\dfrac{\Sigma f.x}{\Sigma f}=\dfrac{0\times8+1\times10+2\times2+3\times5+4\times3+5\times7 }{8+10+2+5+3+7}"

"=\dfrac{10+4+15+12+35}{35}=\dfrac{76}{35}=2.17"